OpenStudy (anonymous):

5 sqrt10(18 - sqrt6)

OpenStudy (anonymous):

Are you asking \[5\sqrt{10(18-\sqrt{6})}\]?

OpenStudy (anonymous):

no 5 sqrt10 (18- sqrt6)

OpenStudy (anonymous):

So... \[5\sqrt{10}(18-\sqrt{6})\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[90\sqrt{10}-5\sqrt{60}\] \[90\sqrt{10}-10\sqrt{15}\]

OpenStudy (anonymous):

how did you get the 15?

OpenStudy (anonymous):

5*sqrt(60) = 5*sqrt(4)*sqrt(15)=5*2*sqrt(15)=10*sqrt(15).

OpenStudy (anonymous):

ok thanks i got it can i ask another question?

OpenStudy (anonymous):

Sure... just post another one, unless your question pertains to the one you just asked me.

OpenStudy (anonymous):

sqrt5 divided by 18 + sqrt3

OpenStudy (anonymous):

Are you trying to rationalize it?

OpenStudy (anonymous):

i don't know. i think i just need to simplify it

OpenStudy (anonymous):

It is basically simplified already. Rationalize means that there can't be any roots in the denominator.

OpenStudy (anonymous):

yes - rationalize it please

OpenStudy (anonymous):

sqrt(5)/(18+sqrt(3)). Multiply both the top and bottom by 18-sqrt(3). On the top, you have sqrt(5)(18-sqrt(3)). On the bottom, you have 18^2-3. I hope you can simplify from here.

OpenStudy (anonymous):

what does the ^ sign mean?

OpenStudy (anonymous):

power or exponentiation.